OpenStudy (anonymous):

Convert (-5, pi/4) to rectangular coordinates

OpenStudy (jdoe0001):

( -5 , pi/4 ) r \(\theta\) keep in mind that \(\bf x = rcos(\theta) \implies (-5)cos\left(\frac{\pi}{4}\right)\\ y = rsin(\theta)\implies (-5)sin\left(\frac{\pi}{4}\right)\)

OpenStudy (jdoe0001):

so now just get the values for \(\bf cos\left(\frac{\pi}{4}\right), sincos\left(\frac{\pi}{4}\right)\) and get your ordered pair :)

OpenStudy (jdoe0001):

\(\bf sin\left(\frac{\pi}{4}\right)\) that is :)